### v1

@@ 4 Band structure perspective on electrochemical interfaces
While it is straightforward to apply electric fields in empirical molecular dynamics deissenbeck2021 , special care must be taken in the context of electronic structure simulations. Fig. 5a shows a schematic representation of a prototypical interface under electric bias between a metal working electrode, explicit electrolyte (H2O) and a computational counter electrode. The counter electrode can be realized as discussed in the preceding section.
@@
In this setup, the electrolyte band gap defines the maximum possible potential drop, that can be applied across the electrolyte: if the applied voltage $U$ equals the band gap $E_{g}$, the energetic position of the conduction band minimum (CBM) at one side becomes equal to the energetic position of the valence band maximum (VBM) at the other side due to band bending. Fig. 5a depicts a situation where the applied voltage $U$ is only slightly less than $E_{g}$. For $U>E_{g}$ dielectric breakthrough sets in. Therefore, the maximum electric field that can be achieved is $\|\vec{E}_{max}\|=E_{g}/d$, where $d$ is the thickness of the electrolyte.
@@
We emphasize that in practical calculations, however, the maximum electric field is often much smaller than $E_{g}/d$: it also depends on the alignment of the Fermi level $E_{F}$ of the working electrode with respect to the band edges of the electrolyte at zero electrode charge (potential of zero charge: PZC). Although $E_{F}$ can be freely adjusted within $E_{g}$, dielectric breakthrough equally occurs if the Fermi level coincides with either the VBM or CBM, respectively, at any location within the simulation cell due to the tilted bands. Such a situation is shown schematically in Fig. 5b: at the PZC, the Fermi level is located in close vicinity to the VBM. Transferring a charge $q$ from the working electrode to the computational electrode - which is possible for the electronic structure shown in Fig. 5a - now lifts the VBM at the right hand side above $E_{F}$, causing dielectric breakthrough. The charge $\Delta q$ contained in those valence states that are lifted above $E_{F}$ flows back into the working electrode, so that the VBM becomes equal to $E_{F}$: as a consequence, the Fermi level is pinned. The maximum critical field $\vec{E}_{max}$ that can be applied is, hence, the one just before Fermi level pinning occurs.
@@
Figure 5: a) Schematic representation of the band structure across an electrode-electrolyte interface. The field is applied via a semiconducting computational counter electrode (CCE) outside the electrolyte region. The Fermi level can be adjusted freely within the electronic gap of the electrolyte and the CCE. b) Sufficiently strong applied fields are able to tilt the bands of the electrolyte, so that the Fermi level straddles either the valence or conduction band edge. Dielectric breakthrough will occur. A corresponding charge transfer between the working electrode and the CCE reduces the field so that the Fermi level is moved back to the band edge. c) Stronger fields can be applied at the same voltage without dielectric breakthrough by moving the counter charge closer to the working electrode surface.
@@
In the present setup, the maximum electric field that can be applied depends on the thickness of the water slab contained within the simulation cell. At constant applied voltage, the total electric field decreases with increasing distance between the working and computational counter electrodes. For desirable cell sizes, these fields often fall short of realistic interfacial electric field strengths. The reason is that nature deposits the counter charges in the form of ions much closer to the electrode surface than possible via the computational counter electrodes depicted in Figs. 5a,b. Ref. deissenbeck2023fse therefore suggested to decouple the position of the computational counter electrode from the edge of the explicit electrolyte region and place the counter charge directly within the electrolyte, cf. Fig. 5c. This allows achieving much stronger electric fields at the same applied voltage without causing dielectric breakthrough.
@@
Here we emphasize that the situation is further complicated by the fact, that the applied electric field $\vec{E}=\varepsilon_{0}^{-1}[\vec{D}-\vec{P}]$ consists of two components: the equal and opposite charges $q$ on the working electrode and the computational counter electrode give rise to an electric displacement $\vec{D}$. In response to the electric displacement, the electrolyte becomes polarized with an ensuing polarization density $\vec{P}$. It is, however, their sum - the total field $\vec{E}$ which is responsible for the band bending illustrated in Fig. 5. Accordingly, one may choose to work within different electrostatic boundary conditions of $\vec{D}=\text{const.}$. or $\vec{E}=\text{const.}$., as outlined in section 2.
@@
As a consequence, it may be tempting to use $\vec{D}=\text{const.}$. and directly put the targeted charge on the working electrode. Unfortunately, the electrolyte is unable to reorganize instantaneously. For strongly polar solvents like water, in equilibrium $\vec{D}$ and $\vec{P}$ cancel out on average to a large degree, due to the large dielectric constant of water. The polarization density in the electrolyte is formed due to molecular reorientation and ion diffusion. These proceed on rather long time scales compared to those accessible in electronic structure simulations. Hence, polarizing a metal-electrolyte interface usually requires several tens of picoseconds deissenbeck2023 . Directly charging the working electrode under $\vec{D}=\text{const.}$. to its target value, without giving the electrolyte sufficient time to polarize, immediately causes dielectric breakthrough, preventing in most cases even convergence of the DFT self-consistency loop.
@@
One may argue that the $\vec{D}=\text{const.}$. approach may be applied safely by increasing the electrode charge $q$ only in sufficiently small increments, cf., e.g., Ref. surendralal2018 . This is true in principle. However, as discussed above, the electrons and ions undergo thermal fluctuations. By extension, their thermal movement creates fluctuations in the polarization density $\vec{P}(t)$, which lead to potential fluctuations on the order of $\sqrt{k_{B}T/C}\sim$ 1 eV. Even if $q$ is changed in only small increments, one has to check carefully and constantly that the polarization density fluctuations do not lead to electric fields $\vec{E}(t)=\varepsilon_{0}^{-1}[\vec{D}-\vec{P}(t)]>\vec{E}_{max}$.
@@
Moreover, the system is driven out of thermodynamic equilibrium every time $q$ is changed incrementally. Changing $q$ too quickly may not only cause dielectric breakthrough, but can also lead to simulation artefacts such as the flying-ice-cube-effect deissenbeck2023 . For all these reasons, boundary conditions of constant electric field $\vec{E}=\text{const.}$. are often preferred. As discussed in section 2, however, they are still unable to describe the correct thermodynamics under reactive conditions. Instead, it is necessary to go beyond $\vec{E}=\text{const.}$. towards fully open thermodynamic boundary conditions.
5 Open thermodynamic boundary conditions
The computational counter electrodes discussed in sections 3 and 4 enable AIMD simulations at finite applied electric field for supercells that are charge-neutral in total, but with a varying number of electrons $n_{e^{-}}$ per supercell and thus, by extension, a varying charge on the working electrode. Based on these concepts, we now turn to discuss the realization of open thermodynamic boundary conditions with respect to $n_{e^{-}}$, as outlined in section 2.
Figure 6: Schematic representation of a DFT supercell under potential control, embedded in its larger external environment. Due to the thermal motion of electrons and ions both inside and outside the supercell, the local Fermi level as well as the surface dipole are fluctuating. By extension, also the system potential $\Phi(t)$ inside the supercell is fluctuating. The simulation system is surrounded by its external environment, which acts as a reservoir of charge and connects the supercell to the macroscopic voltage source, held at the external potential $\Phi_{macro}$, which is the independent thermodynamic variable controlled in experiments. The supercell is open to an exchange of charge driven by potential fluctuations and Faradaic currents. Therefore, an equivalent circuit can be introduced, consisting of a surface resistance $R$, a voltage source $\Phi_{0}$ and a computational electrode, that reproduces inside the DFT supercell the charge dynamics and potential distribution of the full system under reactive conditions.
Any such control scheme relies on a conjugate pair of extensive and intensive quantities - such as entropy and temperature (thermostat), or volume and pressure (barostat) - where the extensive quantity is used to drive the fluctuating intensive one to a desired target value [^11] We note that in the case of thermostats, the entropy is difficult to control in simulations. There is a need, hence, to perform a Legendre transformation in order to control the temperature via the particle kinetic energies instead. For barostats, in contrast, the volume is straightforwardly adjusted. In that respect, the construction of a potentiostat is conceptually much closer to a barostat than to a thermostat.. In the context of potential control, it is straightforwardly shown that the electrode charge $n$ and the electrode potential $\Phi$, form such a pair of conjugate quantities deissenbeck2021 . Since the electrode is connected to an external reservoir of charge, constructing a potentiostat amounts to deriving an equation of motion for the extensive charge $n_{e^{\text{-}}}(t)$, i. e. the exchange of charge between the electrode at potential $\Phi(t)$ and the external reservoir held at $\Phi_{0}$.
This task is complicated by the fact that the potentiostat is intended to describe the exchange of charge with a reservoir not just in static calculations, but in the context of molecular dynamics at finite temperature. The thermal motion of the electrons and ions, as well as the exchange of charge between the electrode and the reservoir, create thermal fluctuations in the electric field and - by extension - in the system potential $\Phi(t)$. In the physical system, the identical argument applies to the external environment directly adjacent to the microscopic region targeted by our simulations. Therefore, our supercell is embedded in the fluctuating electric field created by the thermal motion of the electrons and ions in its local external environment.
It is the role of the potentiostat to mimic the contact of the supercell with this external environment, cf. Fig. 6. If the reservoir of charge were to be held at an exactly constant potential without any thermal fluctuations (corresponding to $T=0$ K), the supercell would be able to dissipate its thermal potential fluctuations into the external reservoir, thereby violating the 2nd law of thermodynamics. Such an artificial net energy drain would constantly drive the simulation system out of equilibrium. In order to prevent such a spurious energy transfer, the electric field fluctuations generated along the MD trajectory must be in thermodynamic equilibrium with the field fluctuations of the external environment at the boundary of the supercell. Therefore, the external reservoir itself must have a finite temperature.
The first such approach was introduced in 2012 by Bonnet et al. otani2012PRL , who drove the exchange of charge between electrode and reservoir via a Nose-Hoover algorithm. The Nose-Hoover approach is known to perform well for reasonably ergodic degrees of freedom. The electrode charge, however, is only a single degree of freedom and thus requires Nose-Hoover chains at the cost of an increased number of parameters (Nose-masses) and associated extra tuning. Moreover, the absence of suitable computational counter electrodes in 2012 led Bonnet et al. to construct their potentiostat for systems carrying a net excess charge. Hence, the Nose-Hoover potentiostat requires the derivative of the total energy with respect to the net excess charge as an input quantity. This quantity is notoriously difficult to obtain in DFT simulations. Those requirements proved to be serious obstacles in implementing the Nose-Hoover potentiostat into existing DFT codes.
In the context of potential control, driving the system potential $\Phi(t)$ towards the external potential $\Phi_{0}$ is the dissipation. Formally, in an equivalent circuit approach, the dissipation can be represented as a resistance $R$ between the part of the working electrode located inside the supercell and its surrounding external environment, cf. Fig. 6. Physically, $R$ describes the surface conductivity of the electrode surface. This dissipation must then be balanced by the energy gained from the thermal potential fluctuations of the local external environment at the boundary of the supercell. In the equivalent circuit approach, the fluctuations are then given by the Johnson-Nyquist noise generated by the resistance $R$. It was shown in Ref. deissenbeck2021 , that the fluctuation-dissipation theorem in this case takes the form
where $f$ is the rate of change of the electrode potential $\Phi$, $\tau_{\Phi}$ is a relaxation time constant and $dW_{t}$ is the time differential of the Wiener stochastic process.
Such a thermopotentiostat was recently derived by Deißenbeck et al. deissenbeck2021 ; deissenbeck2023 via Ito integration gardiner of the fluctuation-dissipation relation shown above. Thereby, the thermopotentiostat provides a direct expression for the charge $n_{e^{-}}$ at each discrete MD time step, under the open thermodynamic boundary conditions depicted in Fig. 6. In contrast to the $D=\text{const.}$. and $E=\text{const.}$. approaches, no artificial constraints on either the charge or potential dynamics are imposed  [^12] In the limit of setting the equivalent resistance $R$ (or by extension the relaxation time $\tau_{\Phi}$) to either zero or infinity, respectively, electrostatic boundary conditions of $\vec{E}=\text{const.}$. or $\vec{D}=\text{const.}$. are recovered.. Instead, the thermopotentiostat is mathematically guaranteed to sample the charge and potential distributions depicted in Fig. 2c, which preserve the reactive area under charge transfer reactions. In addition, it no longer requires the derivative of the total energy with respect to the charge as an input quantity. Instead, $\Phi(t)-\Phi_{0}$ is the only mandatory input quantity. In computational setups such as the ones shown in Figs. 4b, c, e, f this potential difference can be conveniently measured via the total dipole moment - given, e. g., by the dipole correction neugebauer1992adsorbate - along the normal direction of the electrode surface. Therefore, in conjunction with any of those computational setups the thermopotentiostat can be straightforwardly included in any electronic structure package.
@@ 6 Conclusion
Electrochemical interfaces are thermodynamically open systems, that exchange energy, electronic charge as well as ions with their environment. In addition, on the length and time scale relevant for chemical reactions large spatial and temporal fluctuations of the electrostatic potential and field occur. Including these features in realistic ab initio simulations is challenging.
In this review, we have systematically analyzed the fundamental nature and the physics behind these challenges. They are related to three main issues: First, electrostatic phenomena arising from the long-range nature of the Coulomb interaction in small supercells with periodic boundary conditions. Issues discussed here relate to the choice and arrangement of the computational counter electrode. A second main point addressed the phenomena and corrections necessary to describe these interfaces on the small length scale characteristic to DFT supercells. An important point here are the charge and field fluctuations and how they can be consistently included by a careful choice of boundary conditions. The corresponding stochastic potentiostat features the same ease of application as the more commonly used thermostats. Finally, we discussed issues related to band structure and band alignment, which need to be addressed to avoid unphysical simulation artifacts, e.g. by a dielectric breakthrough or to understand limitations in the achievable maximum fields or applied voltages.
Based on this analysis we discussed the available techniques and simulation approaches to address and overcome these limitations. While these methodological developments provide researchers with a new level of realism to study fundamental electrochemical mechanisms and reactions using ab initio techniques, critical aspects remain open.
@@
On one side, these simulations are computationally extremely expensive, limiting fully ab initio MD simulations with explicit water and electric fields to a few dozen picoseconds. This often limits our ability to get thermodynamically well-converged properties, explore rare events such as chemical reactions with high barriers, or to use improved DFT exchange-correlation functionals, e.g. hybrid functionals, which could help to overcome issues related to the severely underestimated bandgap of water when using semi-local functionals such as GGA. The recent success of machine learning potentials shows great potential to substantially speed up such calculations. However, in contrast to many of the existing approaches where short-range potentials with an interaction cutoff radius of a few Angstrom are sufficient, the inherent electrostatic nature of electrochemical interfaces makes the inclusion of the long-range Coulomb interactions mandatory. Alternatively, improved implicit solvation or microsolvation techniques that include a limited amount of explicit solvent directly at the electrified interface are conceivable. Also hybrid approaches, where machine learning is used to describe the dielectric response of the electrolyte to the electrified interface may be promising to reach realistic time scales. Moreover, the computational demand can be reduced even further by advanced sampling techniques, such as, e.g., umbrella sampling or metadynamics laio2002 , upsampling Jung2023 , as well as accelerated molecular dynamics approaches, such as forward flux sampling Allen2006 ; Allen2009 .

### v2

@@ 5 Band structure perspective on electrochemical interfaces
While it is straightforward to apply electric fields in classical molecular dynamics deissenbeck2021 , special care must be taken in the context of electronic structure simulations. Fig. 5a shows a schematic representation of a prototypical interface under electric bias between a metal working electrode, explicit electrolyte (H2O) and a surrogate model for the electric double layer. The surrogate model can be realized as discussed in the preceding section.
@@
In any such setup, the electrolyte band gap defines the maximum possible potential drop, that can be applied across the electrolyte: if the applied voltage $U$ equals the band gap $E_{g}$, the energetic position of the conduction band minimum (CBM) at one side becomes equal to the energetic position of the valence band maximum (VBM) at the other side due to band bending. Fig. 5a depicts a situation where the applied voltage $U$ is only slightly less than $E_{g}$. For $U>E_{g}$, the VBM on the right hand side would be located energetically above the CBM on the left hand side, and, in turn, also above the Fermi energy. Hence, dielectric breakdown sets in, inducing a charge transfer $\delta q$, until the VBM shifts again below the Fermi energy. Therefore, the maximum electric field that can be achieved is $\|\vec{E}_{max}\|=E_{g}/d$, where $d$ is the thickness of the electrolyte.
@@
We emphasize that in practical calculations, however, the maximum electric field that can be achieved before breakdown sets in is often much smaller than $E_{g}/d$: it also depends on the alignment of the Fermi level $E_{F}$ of the working electrode with respect to the band edges of the electrolyte at zero electrode charge (potential of zero charge: PZC). Although $E_{F}$ can be freely adjusted within $E_{g}$, dielectric breakdown equally occurs if the Fermi level coincides with either the VBM or CBM, respectively, at any location within the simulation cell due to the tilted bands. Such a situation is shown schematically in Fig. 5b: at the PZC, the Fermi level is located in close proximity to the VBM. Transferring a charge $q$ from the working electrode to the computational electrode - which is possible for the electronic structure shown in Fig. 5a - now lifts the VBM at the right hand side above $E_{F}$, causing dielectric breakdown. The charge $\Delta q$ contained in those valence states that are lifted above $E_{F}$ flows back into the working electrode, so that the VBM becomes equal to $E_{F}$: as a consequence, the Fermi level is pinned. The maximum critical field $\vec{E}_{max}$ that can be applied is, hence, the one just before Fermi level pinning occurs.
@@
Figure 5: a) Schematic representation of the band structure across an electrode-electrolyte interface. The field is applied via a semiconducting computational counter electrode (CCE) outside the electrolyte region. The Fermi level can be adjusted freely within the electronic gap of the electrolyte and the CCE. b) Sufficiently strong applied fields are able to tilt the bands of the electrolyte, so that the Fermi level straddles either the valence or conduction band edge. Dielectric breakdown will occur. A corresponding charge transfer between the working electrode and the CCE reduces the field so that the Fermi level is moved back to the band edge. c) Stronger fields can be applied at the same voltage without dielectric breakdown by moving the counter charge closer to the working electrode surface.
@@
In any setup with explicit water, the maximum electric field that can be applied depends on the thickness of the water slab contained within the simulation cell. At constant applied voltage, the total electric field decreases with increasing distance between the working and computational counter electrodes. For desirable cell sizes, these fields often fall short of realistic interfacial electric field strengths. The reason is that nature deposits the counter charges in the form of ions much closer to the electrode surface than possible via the computational counter electrodes depicted in Figs. 5a,b. Ref. deissenbeck2023 therefore suggested to decouple the position of the computational counter electrode from the edge of the explicit electrolyte region and place the counter charge directly within the electrolyte, cf. Fig. 5c. This allows achieving much stronger electric fields at the same applied voltage without causing dielectric breakdown.
@@
Here we emphasize that the situation is further complicated by the fact, that the applied electric field $\vec{E}=\varepsilon_{0}^{-1}[\vec{D}-\vec{P}]$ consists of two components: the equal and opposite charges $q$ on the working electrode and the computational counter electrode give rise to an electric displacement $\vec{D}$. In response to the electric displacement, the electrolyte becomes polarized with an ensuing polarization density $\vec{P}$. It is, however, their sum - the total field $\vec{E}$ - which is responsible for the band bending illustrated in Fig. 5. Accordingly, one may choose to work within different electrostatic boundary conditions of $\vec{D}=\text{const.}$ or $\vec{E}=\text{const.}$, as outlined in section 3.
@@
As a consequence, it may be tempting to use $\vec{D}=\text{const.}$ and directly put the targeted charge on the working electrode. Unfortunately, the electrolyte is unable to reorganize instantaneously. For strongly polar solvents like water, in equilibrium $\vec{D}$ and $\vec{P}$ cancel out on average to a large degree, due to the large dielectric constant of water. The polarization density in the electrolyte is formed due to molecular reorientation and ion diffusion. These proceed on rather long time scales compared to those accessible in electronic structure simulations. Hence, polarizing a metal-electrolyte interface usually requires several tens of picoseconds deissenbeck2023 . Directly charging the working electrode under $\vec{D}=\text{const.}$ to its target value, without giving the electrolyte sufficient time to polarize, immediately causes dielectric breakdown, preventing in most cases even convergence of the DFT self-consistency loop.
@@
One may argue that the $\vec{D}=\text{const.}$ approach may be applied safely by increasing the electrode charge $q$ only in sufficiently small increments, cf., e.g., Ref. surendralal2018 . This is true in principle. However, as discussed above, the electrons and ions undergo thermal fluctuations. By extension, their thermal movement creates fluctuations in the polarization density $\vec{P}(t)$, which lead to potential fluctuations on the order of $\sqrt{k_{B}T/C}\sim$ 1 V, where $C$ is the capacitance of the system. Even if $q$ is changed in only small increments, one has to check carefully and constantly that the polarization density fluctuations do not lead to electric fields $\vec{E}(t)=\varepsilon_{0}^{-1}[\vec{D}-\vec{P}(t)]>\vec{E}_{max}$.
@@
Moreover, the system is driven out of thermodynamic equilibrium every time $q$ is changed incrementally. Changing $q$ too quickly may not only cause dielectric breakdown, but can also lead to simulation artefacts such as the flying-ice-cube-effect deissenbeck2023 . For all these reasons, boundary conditions of constant electric field $\vec{E}=\text{const.}$ are often preferred. As discussed in section 3, however, they are still unable to describe the correct thermodynamics under reactive conditions. Instead, it is preferable to go beyond $\vec{E}=\text{const.}$ and use fully open thermodynamic boundary conditions (Sec. 4).
@@ 6 Conclusion
Electrochemical interfaces present a unique challenge for first-principles simulations, as they involve thermodynamically open systems that exchange energy, charge, and ions with their environment. These interfaces are governed by distinct timescales: the rapid charge equilibration of the electrode, driven by electron mobility; the slower ionic motion that defines the double-layer relaxation; and the intermediate timescales associated with electrochemical reactions. The interplay of these timescales leads to dynamic fluctuations in the local potential and electric field, which emerge from the complex, time-dependent interactions between electrode charge, solvent dynamics, and ionic distributions within the double layer.
These fluctuations are particularly relevant for electrochemical reactions, which are inherently local in time and space and therefore strongly influenced by transient potential fluctuations. Capturing these fluctuations is essential not only to replicate the correct mean potential but also to reproduce the thermal equilibrium behavior of the field. The challenge lies in ensuring that the finite DFT supercell reproduces the statistical and dynamic behavior of the effectively infinite electrochemical environment. Achieving this requires electrostatic boundary conditions that allow for the correct distribution of potential fluctuations, ensuring that reaction barriers, field distributions, and interfacial structures are realistically modeled—a task that remains challenging for conventional electronic structure methods.
In this review, we have analyzed the fundamental principles and available techniques for simulating electrochemical interfaces from first principles. We discussed the necessity of surrogate models for the electrochemical double layer and examined how different electrostatic boundary conditions influence potential distributions, reaction rates, and interfacial properties. We emphasized the importance of capturing not only the mean potential but also the fluctuations that characterize the system’s thermal equilibrium, drawing analogies to the well-established role of thermostats in molecular dynamics.
A central finding is that boundary conditions enforcing either constant electric field or constant charge each capture only part of the interfacial physics. Constant field conditions reproduce fast electrode charge dynamics, while constant charge conditions better approximate the sluggish ionic relaxation of the double layer. Thermodynamically open boundary conditions provide a more complete framework by capturing both processes simultaneously. This approach enables simulations at constant potential on average - consistent with experimental conditions - while preserving the natural fluctuations and dynamics that govern reaction barriers and rates. Finally, we discussed issues related to band structure and band alignment, which need to be addressed to avoid unphysical simulation artifacts, e.g. by a dielectric breakdown or to understand limitations in the achievable maximum fields or applied voltages.
While these methodological developments provide researchers with a new level of realism to study fundamental electrochemical mechanisms and reactions using ab initio techniques, critical aspects remain open. An open conceptual challenge remains in the treatment of lateral charge dynamics. While the thermopotentiostat enables controlled charge exchange normal to the electrode surface, lateral directions typically assume instantaneous electron equilibration through periodic boundary conditions – effectively enforcing constant electric field boundary conditions. This reflects the Born-Oppenheimer approximation, which neglects the finite lateral conductivity and surface resistance of real electrodes. In reality, lateral charge redistribution is governed by electron dynamics and finite surface mobility. Recent studies, for example in the context of surface contaminants in ion traps, have shown that finite surface resistance can give rise to lateral fluctuation-induced currents, which in turn affect the motion of nearby ions henkel16 . Such effects may become relevant in systems with spatially extended redox processes or fluctuating surface inhomogeneities. Incorporating explicit electron dynamics – such as real-time TDDFT or Car-Parrinello-like approaches – offers a promising route to go beyond the Born-Oppenheimer approximation and capture lateral equilibration processes, thereby enhancing the realism of potential-controlled simulations.
@@
Moreover, these simulations are computationally extremely expensive, limiting fully ab initio MD simulations with explicit water and electric fields to a few dozen picoseconds. This often limits our ability to get thermodynamically well-converged properties, explore rare events such as chemical reactions with high barriers, or to use improved DFT exchange-correlation functionals, e.g. hybrid functionals, which could help to overcome issues related to the severely underestimated bandgap of water when using semi-local functionals such as GGA. The recent success of machine learning potentials shows great potential to substantially speed up such calculations. However, in contrast to many of the existing approaches where short-range potentials with an interaction cutoff radius of a few Angstrom are sufficient, the inherent electrostatic nature of electrochemical interfaces makes the inclusion of the long-range Coulomb interactions mandatory dellago2023 . Alternatively, improved implicit solvation or microsolvation techniques plaisance2023 that include a limited amount of explicit solvent directly at the electrified interface are conceivable. Hybrid approaches, where machine learning or atomistic calculations ringe2022 are used to describe the dielectric response of the electrolyte to the electrified interface may also be promising to reach realistic time scales. Moreover, the computational demand can be reduced even further by advanced sampling techniques, such as, e.g., umbrella sampling or metadynamics metadyn1 , upsampling Jung2023 , as well as accelerated molecular dynamics approaches, such as forward flux sampling Allen2006 ; Allen2009 .
@@
[…]
Acknowledgements
The authors acknowledge the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for funding through Projects No. 409476157 (SFB1394), No. 506711657 (SFB1625), No. 223848855 (SFB1083) and support under Germany’s Excellence Strategy - EXC 2033-390677874–RESOLV.
Declarations
  • Funding:
    DFG Project No. 409476157 (SFB1394),
    DFG Project No. 506711657 (SFB1625),
    DFG Project No. 223848855 (SFB1083).
  • Conflict of interest/Competing interests: There are no conflicts of interest or competing interests.
  • Ethics approval and consent to participate: Not applicable.
  • Consent for publication: Not applicable.
  • Data availability: Not applicable.
  • Materials availability: Not applicable.
  • Code availability: Not applicable.
  • Author contribution: All authors contributed equally to this work.